OpenStudy (anonymous):

simplify sqrt 216k^4

jimthompson5910 (jim_thompson5910):

Hint: \[\Large \sqrt{216} = \sqrt{36*6}\] \[\Large \sqrt{216} = \sqrt{36}*\sqrt{6}\]